600g of CaCOg decomposes in 2 minutes, what is The rate of reaction


1
Expert's answer
2021-06-21T06:47:08-0400

Rate = Mol/min

Molar Mass of CaCo = 100

= 600/100 = 6 mols

6/2

= 3 Mols/minute